10. MULTIPLE PRODUCT BREAKEVEN POINT. Angus Manufacturing makes basic tools for sale in hardware stores across the country. It is a successful business, with the following unit sales in a month. It faces fixed costs of $225,000 per month. Angus wants you to help him understand how his product mix affects his business. Angus Manufacturing Price per Unit Variable Cost per Unit Product Units Sold Hammers Socket wrenches Plain screwdriver Phillips-head screwdriver Pliers $10 $7.50 25,000 $25 $15.00 10,000 $8 $6.40 10,000 $8 $6.40 5,000 $10 $7.50 25,000 REQUIRED: a. What is Anguss product mix? Base this on Phillips-head screwdrivers. b. What is Anguss package contribution margin? C. Perform a multiproduct breakeven for Angus Manufacturing. How many of each type of product does he need to sell in order to break even? d. What is Anguss current profit for the business? e. What is Anguss current margin of safety in terms of revenue dollars? f. Angus wants to make $100,000 per month before taxes. How many units of each type of product does the company need to sell now? What is its required reve- nue dollars of sales?
Answer #1

a)Product mix based on philips head screwdrivers is
head screwdriver:socket:plain screwdriver:hammers:pliers
=5000:10000:10000:25000:25000
=1:2:2:5:5

b)
package contribution margin=sum(contribution margin of each per unit*product sold)

CM=price per unit-variable cost per unit
Hammers=10-7.5=2.5
Socket=25-15=10
Plain screwdriver=8-6.4=1.6
head screwdriver=8-6/4=1.6
pliers=10-7.5=2.5
package contribution margin=(25000*2.5)+(10000*10)+(10000*1.6)+(5000*1.6)+(25000*2.5)=249000

c)Weighted average contribution margin per unit=total contribution margin/total units
=249000/(25000+10000+10000+5000+25000)=3.32
Breakeven point for entire=fixed costs/cm per unit
=225000/3.32=67771(rounded)
Head screw drivers=(1/(1+2+2+5+5))*67771=4518
Soclet=(2/(1+2+2+5+5))*67771=9036
Plain screw driver=(2/(1+2+2+5+5))*67771=9036
Hammers=(5/(1+2+2+5+5))*67771=22590
pliers=(5/(1+2+2+5+5))*67771=22590

d)Current profit= package conribtuin margin-fixed costs
=249000-225000=24000

e)margin of safety=current sales-breakeven sales
current sales=(25000*10)+(10000*25)+(10000*8)+(5000*8)+(25000*10)=870000
weighted average contribution margin ratio=package contibution margin/sales
=249000/870000=28.62%
sales=225000/28.62%=786145
margin of safety=870000-78614=83855